OpenStudy (anonymous):

What is the 10th term of the geometric sequence –1, 3, –9, …? 19,683 2,187 –59,049 –6,561

OpenStudy (anonymous):

19683

OpenStudy (anonymous):

The equation is just multiply by three, every other number you add a negative sign.

OpenStudy (jamesj):

In the language of geometric series, the series is s1 = a s2 = ar s3 = ar^2 ... sn = ar^(n-1) What is a and r here? Once you've found them, the 10th term, s(10) = ar^9